### v1

# An Analysis of Causal Effect Estimation using Outcome Invariant Data Augmentation

Abstract
The technique of data augmentation (DA) is often used in machine learning for regularization purposes to better generalize under i.i.d. settings. In this work, we present a unifying framework with topics in causal inference to make a case for the use of DA beyond just the i.i.d. setting, but for generalization across interventions as well. Specifically, we argue that when the outcome generating mechanism is invariant to our choice of DA, then such augmentations can effectively be thought of as interventions on the treatment generating mechanism itself. This can potentially help to reduce bias in causal effect estimation arising from hidden confounders. In the presence of such unobserved confounding we typically make use of instrumental variables (IVs)—sources of treatment randomization that are conditionally independent of the outcome. However, IVs may not be as readily available as DA for many applications, which is the main motivation behind this work. By appropriately regularizing IV based estimators, we introduce the concept of IV-like (IVL) regression for mitigating confounding bias and improving predictive performance across interventions even when certain IV properties are relaxed. Finally, we cast parameterized DA as an IVL regression problem and show that when used in composition can simulate a worst-case application of such DA, further improving performance on causal estimation and generalization tasks beyond what simple DA may offer. This is shown both theoretically for the population case and via simulation experiments for the finite sample case using a simple linear example. We also present real data experiments to support our case.
1 Introduction
A classical problem in machine learning is that of regression—using i.i.d. samples from some fixed, unknown distribution $\mathbb{P}_{{X,Y}}$, we predict outcome $Y$ values for unlabelled treatment $X$ values. The use of regularization techniques is crucial for this task to achieve good generalization from training to test data [vapnik]. Data augmentation (DA) [da-reg, invariance_benefit] is one such method, where each sample is randomly perturbed multiple times to grow the dataset size. However, these regression models cannot generally be interpreted causally as the statistical relationship between $X$ and $Y$ may arise from shared common causes, known as confounders, rather than from $X$ influencing $Y$. Removing such confounders requires independently assigning values of $X$ during data generation, known as an intervention [eoci, Pearl_2009].
Unfortunately, we seldom have access to the data generation process to be able to intervene on variables. A common workaround is to use auxiliary variables to correct for unobserved confounders [backdoor, specIV, prox-causal-learning]. One such approach is that of instrumental variables (IVs) that represent certain conditional independences in the system which can be used to identify the causal effect of $X$ on $Y$ [iv_arthur, iv_rui, iv_niki]. Alas, IVs too are generally hard to find in may popular applications such as computer vision and natural language processing, motivating the need for more accessible ways to mitigate unobserved confounding.
Recent work therefore seeks to leverage more commonly available auxiliary variables to reduce confounding-induced bias even when the causal effect itself cannot be identified [causal-regularization2, deconfounding-causal-regularization, regularizing-towards-causal-invariance, anchor]. Collectively referred to as causal regularization, these methods aim to learn predictors that generalize out-of-distribution (OOD) by discouraging reliance on spurious (i.e., non-causal,) correlations. Since distribution shifts often correspond to interventions on parts of the data-generating process [icp, eoci], models that fail under such shifts typically do so because they exploit confounded relationships [when-shift-happens]. Tackling this root cause directly, causal regularization offers a principled approach for more robust prediction.
In the same vein, more ambitious works have also explored the use of common regularization techniques, such as $\ell_{1}$, $\ell_{2}$ [causal-regularization] and the min-norm interpolator [causal-interpolation-regularization], for the same purpose of causal regularization. This is in contrast to the canonical use of such regularizers for estimation variance reduction and i.i.d. prediction generalization [vapnik]. Other popular regularization methods, however, remain understudied in a similar context of un-identifiable causal effect estimation, motivating our work.
Our contributions.
To this end, we provide a first analysis of DA for estimating un-identifiable causal effects using only observational data for $(X,Y)$. Our contributions, summarized in Tab.˜1, include: (i) DA as a soft intervention (Sec.˜4.1): We show that DA can synthesize treatment interventions when the outcome function is invariant to DA, lowering bias in causal effect estimates when the intervention acts along spurious features. (ii) Introducing IV-like regression (Sec.˜3): Relaxing the properties of IVs, we introduce the concept of IV-like (IVL) variables. This generalization renders IV regression ineffective at identifying causal effects, but when regularized appropriately via our proposed IVL regression, may still reduce confounding bias and improve prediction generalization across treatment interventions. (iii) DA parameters as IVL (Sec.˜4.2): By casting parameterized DA as IVL, we show that its composition DA+IVL with IVL regression further reduces confounding bias beyond just simple DA by essentially simulating a worst-case or adversarial application of the DA.
We validate our approach with theoretical results in a linear setting for the infinite-sample case, and simulation and real-data experiments in the finite-sample case.
2 Preliminaries
Consider treatment $X$ and outcome $Y$ taking values in ${\mathcal{X}}\subseteq\mathbb{R}^{m}$ and ${\mathcal{Y}}\subseteq\mathbb{R}^{l}$ respectively. Given the set of functions ${\mathcal{H}}\coloneqq\left\{h:{\mathcal{X}}\rightarrow{\mathcal{Y}}\right\}$, the canonical setting described in the literature [eoci, anchor, econAnalysis] deals with estimating the function $f\in{\mathcal{H}}$ in the structural equation model (SEM) $\mathfrak{M}$ of the following form [^1] Throughout this work we shall borrow and overload notation from [eoci]. See Appendix for a list of symbols.
$X=\tau\left(Y,Z,C,N_{X}\right),$ $Y=f\left(X\right)+\epsilon\left(C\right)+N_{Y},$ (1)
where $Z$, $C$, $N_{X}$, $N_{Y}$ are exogenous (and therefore mutually independent) random variables and the residual $\xi\coloneqq Y-f\left(X\right)=\epsilon\left(C\right)+N_{Y}$ is assumed to be zero mean, i.e. $\mathbb{E}^{\mathfrak{M}}\left[\xi\right]=0$. Since $\mathfrak{M}$ is potentially cyclic, a priori it may entail several or no distributions at all. However, here we make the assumption that for all $\left({\mathbf{x}}_{0},{\mathbf{y}}_{0}\right)\in{\mathcal{X}}\times{\mathcal{Y}}$ the unique limits
${\mathbf{x}}\coloneqq\lim_{t\rightarrow\infty}{\mathbf{x}}_{t}=\lim_{t\rightarrow\infty}\tau\left({\mathbf{y}}_{t-1},{\mathbf{z}},{\mathbf{c}},{\mathbf{n}}_{X}\right),$ ${\mathbf{y}}\coloneqq\lim_{t\rightarrow\infty}{\mathbf{y}}_{t}=\lim_{t\rightarrow\infty}f\left({\mathbf{x}}_{t-1}\right)+\epsilon\left({\mathbf{c}}\right)+{\mathbf{n}}_{Y}$
exist for any $\left({\mathbf{z}},{\mathbf{c}},{\mathbf{n}}_{X},{\mathbf{n}}_{Y}\right)\sim\mathbb{P}_{{Z,C,N_{X},N_{Y}}}^{\mathfrak{M}}$, meaning that the unique distribution entailed by $\mathfrak{M}$ is in this equilibrium state. Of course, if $\mathfrak{M}$ is acyclic, these limits always exist. Note that assuming the existence of such an equilibrium does not violate the classic independent causal mechanism (ICM) principle [eoci]; we defer interested readers to LABEL:app:cyclic-sem for further details on cyclic SEMs and the ICM.
Table 1: A picture summary of our contributions. $\rightarrow$ represents composition of operations or transformations, and $\Leftrightarrow$ represents equivalence.
Type of Data Augmentation Topics in Causal Inference
None; observational data $\leftarrow$ Data generating structural model
$\downarrow$ $\downarrow$
[…]
$\downarrow$ $\downarrow$
[…]
Lower confounding bias in causal effect estimate
(a) Graph of $\mathfrak{M}$.
(b) Graph of $\mathfrak{M};\operatorname{do}\left(X\right)$.
Figure 1: Graph of $\mathfrak{M}$ depicting an instrument $Z$ that satisfies treatment relevance, exclusion restriction, un-confoundedness and outcome relevance properties. An intervention on $X$ gives us the graph in (b). IV regression simulates such an intervention using only observational data.
Given a proper convex loss $\ell:\mathbb{R}^{l}\times\mathbb{R}^{l}\rightarrow\mathbb{R}_{+}$, empirical risk minimization (ERM) uses a dataset ${\mathcal{D}}\coloneqq\left\{({\mathbf{x}}_{i},{\mathbf{y}}_{i})\right\}_{i=0}^{n}$ of $n$ samples from $\mathfrak{M}$ to minimize an empirical version of the statistical risk
$R_{\text{ERM}}^{\mathfrak{M}}\left(h\right)\coloneqq\mathbb{E}^{\mathfrak{M}}\left[\ell\left(Y,h\left(X\right)\right)\right],$ (2)
over $h\in{\mathcal{H}}$. However, since the residual $\xi$ in Eq.˜1 is generally correlated with $X$, i.e., $\mathbb{E}^{\mathfrak{M}}\left[\xi\;\middle|\;X\right]\neq 0$, the ERM minimizer $\hat{h}^{\mathfrak{M}}_{\text{ERM}}$ typically yields a biased estimate of $f$ [Pearl_2009, eoci]. This bias arises due to the exclusion of the (unobserved) common parent $C$ of $X$ and $Y$, i.e. a confounder, in the ERM objective (hence fittingly called the omitted-variable bias [omitted-variable-bias]) and/or the model is cyclic (simultaneity bias [econAnalysis, cyclic-sem-iv], or reverse causality [Pearl_2009] in the degenerate case). For simplicity we shall refer to either case by saying that $X$ and $Y$ are confounded and the resulting bias as the confounding bias [Pearl_2009]. [^2] Pearl [Pearl_2009, p.78,184] similarly uses the term for any bias causing observational vs. interventional deviation; this also aligns with econometrics [trinity, econAnalysis], where both are classified as sources of endogeneity (i.e., $X\not\!\perp\!\!\!\perp\xi$).
2.1 Intervention for causal effect estimation
We can make $X$ and the residual $\xi$ uncorrelated via an intervention [^3] A soft intervention replaces the mechanism $\tau$ in Eq. 1 with an alternative $\tau^{\prime}$ [eoci, p. 34]. This may potentially reduce confounding between $X$ and $Y$. $\operatorname{do}\left(X\coloneqq X^{\prime}\right)$, where we explicitly set $X$ to some independently sampled $X^{\prime}$ in Eq.˜1 irrespective of its parents, resulting now in the new SEM $\mathfrak{M};\operatorname{do}\left(X\coloneqq X^{\prime}\right)$ or $\mathfrak{M};\operatorname{do}\left(X\right)$ as a shorthand for when $X^{\prime}\sim\mathbb{P}_{{X}}^{\mathfrak{M}}$. The distribution induced by this modified SEM is called an interventional distribution (with respect to $\mathfrak{M}$) under which the ERM objective from Eq.˜2 now defines the following causal risk (CR) [causal-regularization2, causal-interpolation-regularization, optical-device] as
$R_{\text{CR}}^{\mathfrak{M}}\left(h\right)\coloneqq R_{\text{ERM}}^{\mathfrak{M};\operatorname{do}\left(X\right)}\left(h\right)=R_{\text{ERM}}^{\mathfrak{M};\operatorname{do}\left(X\coloneqq X^{\prime}\right)}\left(h\right),$ s.t. $X^{\prime}\sim\mathbb{P}_{{X}}^{\mathfrak{M}}.$ (3)
Minimizing Eq.˜3 is meaningful in two important cases where ERM fails: (i) Causal effect estimation: The minimizer $\hat{h}^{\mathfrak{M}}_{\text{CR}}$ of Eq.˜3 gives us an unbiased estimate of the average treatment effect (ATE) [backdoor] $\mathbb{E}^{\mathfrak{M};\operatorname{do}\left(X\coloneqq{\mathbf{x}}\right)}\left[Y\;\middle|\;X={\mathbf{x}}\right]=f\left({\mathbf{x}}\right)$ that measures the causal influence of $X$ on $Y$. (ii) Robust prediction: ATE based prediction of $Y$ values for unlabelled $X$ values is robust in the sense that it can generalize across arbitrary OOD treatment interventions or shifts in the treatment distribution [causal-ood-generalizatoin]. Consequently, the causal risk minimizer $\hat{h}^{\mathfrak{M}}_{\text{CR}}$ is also a robust predictor over the support of $\mathbb{P}_{{X}}^{\mathfrak{M}}$. Specifically, $\hat{h}^{\mathfrak{M}}_{\text{CR}}$ minimizes the worst-case ERM objective over the set $\mathcal{P}$} of all possible intervention distributions $\mathbb{P}_{{X^{\prime}}}$ over the support of $\mathbb{P}_{{X}}^{\mathfrak{M}}$ [causal-ood-generalizatoin], i.e. for ${\mathcal{P}}\coloneqq\left\{\mathbb{P}_{{X^{\prime}}}\;\middle|\;\operatorname{supp}\left(\mathbb{P}_{{X^{\prime}}}\right)\subseteq\operatorname{supp}\left(\mathbb{P}_{{X}}^{\mathfrak{M}}\right)\right\}$,
$\hat{h}^{\mathfrak{M}}_{\text{CR}}\in\operatorname*{argmin}_{h\in{\mathcal{H}}}\max_{\mathbb{P}_{{X^{\prime}}}\in{\mathcal{P}}}R_{\text{ERM}}^{\mathfrak{M};\operatorname{do}\left(X\coloneqq X^{\prime}\right)}\left(h\right).$
To better isolate the estimation error due to confounding, we define the causal excess risk (CER) [causal-interpolation-regularization]
$\operatorname{CER}_{\mathfrak{M}}\left(h\right)\coloneqq R_{\text{CR}}^{\mathfrak{M}}\left(h\right)-R_{\text{CR}}^{\mathfrak{M}}\left(f\right).$
This removes the irreducible noise from Eq.˜3 (see LABEL:app:confounding-bias) and directly measures how far a hypothesis $h$ deviates from the true causal function $f$ under interventions, so that $\operatorname{CER}_{\mathfrak{M}}\left(f\right)=0$.
Since interventions are often inaccessible for computing the risk in Eq.˜3, we usually rely on observational data/ distribution and additional variables to approximate them, as outlined in the next section.
2.2 Instrumental variable regression
One way to get an unbiased estimate of $f$ from the observational distribution of $\mathfrak{M}$ is to use so-called instrumental variables $Z$ with the properties [Pearl_2009, eoci, iv_rui, iv_arthur, dual-iv] of: (i) Treatment Relevance: $Z\not\!\perp\!\!\!\perp X$. (ii) Exclusion Restriction: $Z$ enters $Y$ only through $X$, i.e. $Z\perp\!\!\!\perp Y^{\mathfrak{M};\operatorname{do}\left(X\coloneqq{\mathbf{x}}\right)}$. [^4] Counterfactual definition of the exclusion restriction property [Pearl_2009, p. 248]. (iii) Un-confoundedness: $Z\perp\!\!\!\perp\xi$. (iv) Outcome Relevance: $Z$ carries information about $Y$, i.e. $Y\not\!\perp\!\!\!\perp Z$.
Conditioning Eq.˜1 on $Z$ and using $\mathbb{E}\left[\xi\;\middle|\;Z\right]=\mathbb{E}\left[\xi\right]=0$ from the unconfoundedness property gives
$\mathbb{E}^{\mathfrak{M}}\left[Y\;\middle|\;Z\right]=\mathbb{E}^{\mathfrak{M}}\left[f(X)\;\middle|\;Z\right].$ (4)
IV regression therefore entails solving Eq.˜4 for $f$, which can be done by minimizing the risk [dual-iv]
$R_{\text{IV}}^{\mathfrak{M}}\left(h\right)\coloneqq\mathbb{E}^{\mathfrak{M}}\left[\ell\left(Y,\mathbb{E}^{\mathfrak{M}}\left[h\left(X\right)\;\middle|\;Z\right]\right)\right].$ (5)
For linear $f(\cdot)\coloneqq\mathbf{f}^{\top}(\cdot),h(\cdot)\coloneqq\mathbf{h}^{\top}(\cdot)$ with $\mathbf{f},\mathbf{h}\in\mathbb{R}^{m}$ and squared loss $\ell\left({\mathbf{y}},{\mathbf{y}}^{\prime}\right)\coloneqq\left\lVert{\mathbf{y}}-{\mathbf{y}}^{\prime}\right\rVert^{2}$, this gives the two-stage-least-squares (2SLS) [2SLS-3SLS] solution where the first stage regresses $X$ from $Z$, and the second stage regresses $Y$ from predictions $\mathbb{E}\left[X\;\middle|\;Z\right]$ of the first stage to get the estimate $\hat{h}_{\text{IV}}^{\mathfrak{M}}$.
2.3 Data augmentation
In this work we restrict ourselves to data augmentation with respect to which $f$ is invariant [invariance_benefit, group-theory-da-jmlr]. The action of a group $\mathcal{G}$} is a mapping $\delta:{\mathcal{X}}\times{\mathcal{G}}\rightarrow{\mathcal{X}}$ which is compatible with the group operation. For convenience we shall write ${\mathbf{g}}{\mathbf{x}}\coloneqq\delta({\mathbf{x}},{\mathbf{g}})$. We say that $f$ is invariant under $\mathcal{G}$} (or $\mathcal{G}$}-invariant) if
$f({\mathbf{g}}{\mathbf{x}})=f({\mathbf{x}}),\qquad\forall\;({\mathbf{g}},{\mathbf{x}})\in{\mathcal{G}}\times{\mathcal{X}}.$
Less formally, we say that the map ${\mathbf{g}}{\mathbf{x}}$, henceforth assumed to be continuous in $\mathbf{x}$}, is a valid outcome-invariant DA transformation parameterized by the vector ${\mathbf{g}}\in{\mathcal{G}}$. Let $\mathcal{G}$} have a (unique) normalized Haar measure and $\mathbb{P}_{{G}}$ be the corresponding distribution defined over it. For some $G\sim\mathbb{P}_{{G}}$, the canonical application of DA seeks to minimize an empirical version of the following risk.
$R_{\text{DA}_{G}+\text{ERM}}^{\mathfrak{M}}\left(h\right)\coloneqq\mathbb{E}^{\mathfrak{M}}\left[\ell\left(Y,h\left(GX\right)\right)\right].$ (6)
Note that it is sufficient to have some prior information about the symmetries of $f$ in order to be able to construct such a DA. For example, when classifying images of cats and dogs we already know that whatever the true labeling function may be, it would certainly be invariant to rotations on the images. $G$ would then represent the random rotation angle, whereas $G{\mathbf{x}}$ would be the rotated image $\mathbf{x}$}.
We wish to contrast the use of DA in this work with the canonical setting—to mitigate overfitting, DA is used to grow the sample size by generating multiple augmentations $(G{\mathbf{x}},{\mathbf{y}})$ for each data sample $({\mathbf{x}},{\mathbf{y}})\sim\mathbb{P}_{{{X,Y}}}^{\mathfrak{M}}$ [invariance_benefit, group-theory-da-jmlr, savkin2020adversarial]. Such regularization, overfitting mitigation, estimation variance reduction, or i.i.d. prediction generalization is not the focus of this work and we intentionally provide Eq.˜6 along with theoretical results that follow in the population case to emphasize that DA is not being used as a conventional regularizer. Instead, our goal is to improve causal effect estimation and robust prediction by re-purposing DA to mitigate hidden confounding bias in the data.
3 Faithfulness and Outcome Relevance in IVs
The distribution $\mathbb{P}_{{X,Y,Z,C}}^{\mathfrak{M}}$ is faithful to the graph of $\mathfrak{M}$ if it only exhibits independences implied by the graph [eoci, pgms]. [^5] Also known as stability in some texts [Pearl_2009, p. 48]. This standard assumption in IV settings renders outcome-relevance implicit and therefore rarely mentioned. In this section we discuss the case where only the first three IV properties are satisfied, i.e. outcome-relevance may not hold. Since such a $Z$ may not be a valid IV, therefore identifiability of ATE is not possible in general as the problem in Eq.˜4 can now be misspecified, having multiple, potentially infinitely many solutions when $Y\perp\!\!\!\perp Z$. Nevertheless, we shall refer to such a $Z$ as IV-like (IVL) to emphasize that while $Z$ may not be an IV, it may still be ‘instrumental’ for reducing confounding bias when estimating the ATE compared to the standard ERM baseline.
ERM regularized IV regression.
Despite problem miss-specification for a IVL $Z$, the target function $f$ remains a minimizer for the IV risk in Eq.˜5. Albeit, potentially not unique—for example, a linear $h$ with squared loss leads to an under-determined problem in Eq.˜5. We therefore propose the following regularized version of the IV risk for such an IVL setting,
$R_{\text{IVL}_{\alpha}^{\text{}}}^{\mathfrak{M}}\left(h\right)\coloneqq R_{\text{IV}}^{\mathfrak{M}}\left(h\right)+\alpha R_{\text{ERM}}^{\mathfrak{M}}\left(h\right),$ (7)
where $\alpha>0$ is the regularization parameter. The ERM risk as a penalty allows our estimations to have good predictive performance while the IV risk encourages solution search within the subspace where we know $f$ to be present. We refer to minimising the risk in Eq.˜7 as IVL regression.
Note that the motivation behind IVL regression is not the identifiability of $f$, but rather potentially better estimations of $f$ with lower confounding bias. The next section provides a concrete example.
Example 1 (a linear Gaussian IVL example).
For scalar $\sigma>0$, non-zero matrices ${\bm{\Gamma}},\mathbf{T}\in\mathbb{R}^{\ast\times m}$ and vectors $\bm{\tau}^{\top},\mathbf{f},\bm{\epsilon}\in\mathbb{R}^{m}$ such that $\mathbf{f}^{\top}\bm{\tau}^{\top}\neq 1$ so that the following SEM $\mathfrak{M}$ is solvable in $\left(X,Y\right)$ [^6] See LABEL:app:cyclic-sem and LABEL:lemma:sem-solvability for details on solving for and sampling of $\left(X,Y\right)$ in such linear, cyclic SEMs.
$X=\bm{\tau}^{\top}Y+{\bm{\Gamma}}^{\top}Z+\mathbf{T}^{\top}C+\sigma N_{X},$ $Y=\mathbf{f}^{\top}X+\bm{\epsilon}^{\top}C+\sigma N_{Y},$
where $Z,C,N_{X},N_{Y}$ are conformable, centered Gaussian random vectors and $Z$ is IVL w.r.t. $\left(X,Y\right)$. [^7] All examples assume correlated $X$ and residual $\xi$, i.e. $\mathbb{E}^{\mathfrak{M}}\left[X\xi^{\top}\right]\neq{\mathbf{0}}$, as otherwise there is no confounding.
Now, the task is to improve our estimation of $\mathbf{f}$ compared to standard ERM. We evaluate an estimate $\hat{\mathbf{h}}^{{\mathcal{D}}}$ using the $\operatorname{CER}$, which for a squared loss and covariance ${\bm{\Sigma}}_{X}^{\mathfrak{M}}$ in Example˜1 simply comes out to be
$\operatorname{CER}_{\mathfrak{M}}\left(\hat{\mathbf{h}}^{{\mathcal{D}}}\right)=\left\lVert\hat{\mathbf{h}}^{{\mathcal{D}}}-\mathbf{f}\right\rVert^{2}_{{\bm{\Sigma}}_{X}^{\mathfrak{M}}}.$ (8)
Prior works use this form to quantify the error in ATE estimation [causal-interpolation-regularization, causal-regularization2] or measure some notion of strength of confounding [causal-regularization, dominik-optical-device-causal-reg2, optical-device]. Similarly, we use it to measure confounding bias of population estimates $\hat{\mathbf{h}}^{\mathfrak{M}}$ (LABEL:app:confounding-bias) and estimation error in finite sample experiments. The next results follow.
Theorem 1 (robust prediction with IVL regression).
For SEM $\mathfrak{M}$ in Example˜1, the following holds:
$\hat{\mathbf{h}}^{\mathfrak{M}}_{\text{IVL}_{\alpha}^{\text{}}}\in\operatorname*{argmin}_{\mathbf{h}}\max_{\bm{\zeta}\in\mathcal{P}_{\alpha}}R_{\text{ERM}}^{\mathfrak{M};\operatorname{do}\left({\bm{\Gamma}}^{\top}\left(\cdot\right)\coloneqq\bm{\zeta}\right)}\left(\mathbf{h}\right),$ s.t. $\mathcal{P}_{\alpha}\coloneqq\left\{\bm{\zeta}\;\middle|\;\bm{\zeta}\bm{\zeta}^{\top}\preccurlyeq\left(\frac{1}{\alpha}+1\right){\bm{\Gamma}}^{\top}{\bm{\Sigma}}_{Z}^{\mathfrak{M}}{\bm{\Gamma}}\right\}.$
Proof. See LABEL:app:theorem-ivl-robust-prediction for the proof. ∎
Theorem 2 (causal estimation with IVL regression).
In SEM $\mathfrak{M}$ of Example˜1, for $\alpha<\infty$, we have
$\operatorname{CER}_{\mathfrak{M}}\left(\hat{\mathbf{h}}^{\mathfrak{M}}_{\text{IVL}_{\alpha}^{\text{}}}\right)\leq\operatorname{CER}_{\mathfrak{M}}\left(\hat{\mathbf{h}}^{\mathfrak{M}}_{\text{ERM}}\right),$ equality iff $\mathbb{E}^{\mathfrak{M}}\left[X\;\middle|\;Z\right]\,\perp_{\mathrm{a.s.}}\,\!\mathbb{E}^{\mathfrak{M}}\left[X\;\middle|\;\xi\right].$
Proof. See LABEL:app:theorem-ivl-causal-estimation for the proof. ∎
Theorem˜1 shows that IVL regression achieves optimal predictive performance across treatment interventions within the perturbation set ${\mathcal{P}}_{\alpha}$ defined by $\alpha$. Theorem˜2 further states that this strictly reduces confounding bias in ATE estimates iff the perturbations align with spurious features of $X$, as indicated by the equality condition (also necessary for identifiability in linear IV settings [causal-identifiability-ate, causal-ood-generalizatoin]).
4 Causal Effect Estimation using Data Augmentation
(a) Graph of $\mathfrak{A}$ post DA.
(b) Graph of $\mathfrak{A};\operatorname{do}\left(\tau\coloneqq G\tau\right)$
Figure 1: The observational distribution of $(GX,Y,G,C)$ and $(X,Y,G,C)$ for graphs (a) and (b) respectively are the same. The former applies DA on $X$, whereas the later applies a (soft) intervention on $X$. Furthermore, for the graph in (b), $G$ is IVL.
Figure 2: The ground truth function $\mathbf{f}$ in Example˜2. The DA applied here corresponds to randomly translating the data samples along their level-set by adding random noise sampled from the null-space of $\mathbf{f}$.
We dedicate this section to the main topic and point of this work—discussing the potential of data augmentation for improving predictive performance across interventions and reducing confounding bias in ATE estimates. To that effect, for the rest of this work we shall consider the following SEM $\mathfrak{A}$
$X=\tau\left(Y,C,N_{X}\right),$ $Y=f\left(X\right)+\epsilon\left(C\right)+N_{Y},$ (9)
which is assumed to have a unique stationary distribution with exogenous $C,N_{X},N_{Y}$ and the residual $\xi\coloneqq Y-f\left(X\right)$ is zero-mean, i.e. $\mathbb{E}\left[\xi\right]=0$. We also have access to DA transformations $GX$ of $X$ parameterized by $G\sim\mathbb{P}_{{G}}^{{\mathfrak{A}}}$ such as described in Sec.˜2.3. Figure˜2(a) shows the graph of $\mathfrak{A}$ post DA.
Given samples for only $\left(X,Y\right)$ and some valid DA parameterised by $G$, the task is to improve predictive performance across interventions and reduce confounding bias in ATE estimates. We now make two observations in the following sections and state the respective results that follow thereof.
4.1 Data augmentation as a soft intervention
Consider a (soft) intervention on $\mathfrak{A}$ where we substitute the mechanism $\tau$ of $X$ with $G\tau$. With some abuse of notation, we shall represent this SEM by $\mathfrak{A};\operatorname{do}\left(\tau\coloneqq G\tau\right)$ the graph of which is shown in Fig.˜2(b). Note that this SEM also has a unique stationary distribution (proof in LABEL:app:proposition-da-intervention-distribution-existence). Comparing the DA mechanism in $\mathfrak{A}$ (Fig.˜2(a)) and the intervention $\mathfrak{A};\operatorname{do}\left(\tau\coloneqq G\tau\right)$ (Fig.˜2(b)), we see:
Observation 1 (soft intervention with DA).
Distributions $\mathbb{P}_{{GX,Y,G,C}}^{\mathfrak{A}}$ and $\mathbb{P}_{{X,Y,G,C}}^{\mathfrak{A};\operatorname{do}\left(\tau\coloneqq G\tau\right)}$ are identical.
We can hence treat samples generated from $\mathfrak{A}$ via DA as if they were instead generated from $\mathfrak{A};\operatorname{do}\left(\tau\coloneqq G\tau\right)$ by intervening on $X$. This allows us to re-write the DA+ERM risk from Eq.˜6 as,
$R_{\text{DA}_{G}+\text{ERM}}^{\mathfrak{A}}\left(h\right)=R_{\text{ERM}}^{\mathfrak{A};\operatorname{do}\left(\tau\coloneqq G\tau\right)}\left(h\right),$
to emphasize that DA is equivalent to a (soft) intervention and as such can be used to reduce confounding bias when estimating $f$, as we will show with the following example.
Example 2 (a linear Gaussian DA example).
For scalars $\kappa,\sigma>0$, non-zero matrices ${\bm{\Gamma}},\mathbf{T}\in\mathbb{R}^{\ast\times m}$ and vectors $\bm{\tau}^{\top},\mathbf{f},\bm{\epsilon}\in\mathbb{R}^{m}$ such that $\mathbf{f}^{\top}\bm{\tau}^{\top}\neq\kappa^{-1}$ so that the following SEM $\mathfrak{A}$ is solvable in $(X,Y)$
$X=\kappa\cdot\bm{\tau}^{\top}Y+\mathbf{T}^{\top}C+\sigma N_{X},$ $Y=\mathbf{f}^{\top}X+\kappa\cdot\bm{\epsilon}^{\top}C+\sigma N_{Y},$ $GX\coloneqq X+\gamma\cdot{\bm{\Gamma}}^{\top}G,$
where $G,C,N_{X},N_{Y}$ are conformable, centered Gaussian random vectors, $\kappa$ determines how much $\left(X,Y\right)$ are confounded and $\operatorname{range}\left({\bm{\Gamma}}^{\top}\right)\subseteq\operatorname{null}\left(\mathbf{f}^{\top}\right)$ so that $GX$ is a valid outcome invariant DA transformation of $X$ parameterized by $G$ with strength $\gamma>0$. This transformation can be viewed as translating $X$ along its level-set as shown in Fig.˜2 and represents our prior knowledge about the symmetries of $\mathbf{f}$ for the purposes of this example.
Theorem 3 (causal estimation with DA+ERM).
For SEM $\mathfrak{A}$ in Example˜2, the following holds:
$\operatorname{CER}_{\mathfrak{A}}\left(\hat{\mathbf{h}}^{\mathfrak{A}}_{\text{DA}_{G}+\text{ERM}}\right)\leq\operatorname{CER}_{\mathfrak{A}}\left(\hat{\mathbf{h}}^{\mathfrak{A}}_{\text{ERM}}\right),$ equality iff $\mathbb{E}^{\mathfrak{A}}\left[GX\;\middle|\;G\right]\,\perp_{\mathrm{a.s.}}\,\!\mathbb{E}^{\mathfrak{A}}\left[X\;\middle|\;\xi\right].$
Proof. See LABEL:app:theorem-da-causal-estimation for the proof. ∎
That is, DA strictly reduces confounding bias in ATE estimate iff the induced intervention perturbes $X$ along spurious features. Importantly, Theorem˜3 suggests that lower confounding bias is not a ‘free lunch’ with outcome invariance of DA and practitioners may need domain knowledge to construct DA that targets spurious features. Fortunately however, Theorem˜3 also suggests that with outcome invariance, DA should not perform worse than ERM. We say that DA+ERM dominates ERM on causal estimation [domination, p. 48]. Practitioners may therefore be advised to generously use such DA, as it achieves regularization in the worst case, and mitigates confounding bias as a ‘bonus’ in the best case.
4.2 Worst-case data augmentation with IVL regression
We once again point our attention to the graph of $\mathfrak{A};\operatorname{do}\left(\tau\coloneqq G\tau\right)$ from Fig.˜2(b) to observe that:
Observation 2 (IV-like DA parameters).
In SEM $\mathfrak{A};\operatorname{do}\left(\tau\coloneqq G\tau\right)$, the DA parameters $G$ are IVL.
In light of this we can now re-write the IV and IVL risks for $\mathfrak{A};\operatorname{do}\left(\tau\coloneqq G\tau\right)$ to respectively read
$R_{\text{DA}_{G}+\text{IV}}^{\mathfrak{A}}\left(h\right)=R_{\text{IV}}^{\mathfrak{A};\operatorname{do}\left(\tau\coloneqq G\tau\right)}\left(h\right),$ $R_{\text{DA}_{G}+\text{IVL}_{\alpha}^{\text{}}}^{\mathfrak{A}}\left(h\right)=R_{\text{IVL}_{\alpha}^{\text{}}}^{\mathfrak{A};\operatorname{do}\left(\tau\coloneqq G\tau\right)}\left(h\right).$
Corollary 1 (worst-case DA with DA+IVL regression).
For SEM $\mathfrak{A}$ in Example˜2, it holds that
$\hat{\mathbf{h}}^{\mathfrak{A}}_{\text{DA}_{G}+\text{IVL}_{\alpha}^{\text{}}}\in\operatorname*{argmin}_{\mathbf{h}}\max_{{\mathbf{g}}\in{\mathcal{G}}_{\alpha}}R_{\text{DA}_{{\mathbf{g}}}+\text{ERM}}^{\mathfrak{A}}\left(\mathbf{h}\right),$ s.t. ${\mathcal{G}}_{\alpha}\coloneqq\left\{{\mathbf{g}}\;\middle|\;{\bm{\Gamma}}^{\top}{\mathbf{g}}{\mathbf{g}}^{\top}{\bm{\Gamma}}\preccurlyeq\left(\frac{1}{\alpha}+1\right){\bm{\Gamma}}^{\top}{\bm{\Sigma}}_{G}^{\mathfrak{A}}{\bm{\Gamma}}\right\}.$
Proof. The result follows from ˜1, ˜2 and Theorem˜1. ∎
Corollary 2 (causal estimation with DA+IVL regression).
For $\alpha,\gamma<\infty$ in SEM $\mathfrak{A}$ from Example˜2,
$\operatorname{CER}_{\mathfrak{A}}\left(\hat{\mathbf{h}}^{\mathfrak{A}}_{\text{DA}_{G}+\text{IVL}_{\alpha}^{\text{}}}\right)\leq\operatorname{CER}_{\mathfrak{A}}\left(\hat{\mathbf{h}}^{\mathfrak{A}}_{\text{DA}_{G}+\text{ERM}}\right),$ equality iff $\mathbb{E}^{\mathfrak{A}}\left[GX\;\middle|\;G\right]\,\perp_{\mathrm{a.s.}}\,\!\mathbb{E}^{\mathfrak{A}}\left[X\;\middle|\;\xi\right].$
Proof. The result follows directly from Theorem˜2 and ˜2. ∎
Using DA parameters as IVL therefore simulates a worst-case, or adversarial application of DA within a set of transforms ${\mathcal{G}}_{\alpha}$. Of course Corollary˜1 can also be viewed as a predictor that generalizes to treatment interventions encoded by ${\mathcal{G}}_{\alpha}$. As is intuitive, such a worst-case intervention improves our ATE estimation so long as the features of $X$ intervened along include some that are spurious (Corollary˜2). DA and IVL regression may therefore be used in composition if the application can benefit from regularization and/ or better prediction generalization across DA-induced interventions, with a ‘bonus’ of lower confounding bias if the DA also augments any spurious features of $X$.
5 Related Work
(a) $\gamma=1$, $\kappa\in[0,1]$
(b) $\gamma,\kappa=1$, $\alpha\in[10^{-5},10^{5}]$
(c) $\gamma\in[10^{-2.5},10]$, $\kappa=1$
Figure 3: Simulation experiment for a linear Gaussian SEM. $\kappa$ represents the amount of confounding, $\gamma$ is the strength of DA and $\alpha$ is the IVL regularization parameter. Each data-point represents the average $\operatorname{nCER}$} over 2525 trials with a $95\%$ confidence interval (CI).
Causal regularization is perhaps the most appropriate classification for this work. These methods aim for more robust prediction by mitigating the upstream problem of confounding bias in a more accessible way than is required for full identification. This is done, for example, by relaxing properties of auxiliary variables [causal-regularization2, deconfounding-causal-regularization, regularizing-towards-causal-invariance, anchor], as we have done via our IVL approach. Most relevant, however, are methods that re-purpose common regularizers, canonically used for estimation variance reduction and i.i.d. prediction generalization, for confounding bias mitigation. Of note is [causal-regularization], where a certain linear modeling assumption allows the estimation of $\left\lVert\mathbf{f}\right\rVert^{2}$ from observational $(X,Y)$ data, which is then used to develop a cross-validation scheme for $\ell_{1},\ell_{2}$ regularization. [causal-interpolation-regularization] conducted a similar theoretical analysis for the min-norm interpolator. To the best of our knowledge, we are the first to study the same for DA—re-purposing yet another ubiquitous regularizer to mitigate confounding bias.
Domain generalization (DG) [krikamol-domain-generalization] methods aim for prediction generalization to unseen test domains via robust optimization (RO) [RO] over a perturbation set $\mathcal{P}$} of possible test domains $\rho\in{\mathcal{P}}$ as
$R_{\text{{RO}}}^{{\mathcal{P}}}\left(h\right)\coloneqq\max_{\rho\in{\mathcal{P}}}R_{\text{ERM}}^{\rho}\left(h\right),$
Since generalizing to arbitrary test domains is impossible, the choice of perturbation set encodes one’s assumptions about which test domains might be encountered. Instead of making such assumptions a priori, it is often assumed to have access to data from multiple training domains which can inform one’s choice of perturbation set. This setting is explored in group distributionally robust optimization (DRO) [dro]. Variations have been used to mitigate confounding bias and subsequently generalize to treatment interventions when used with interventional data [icp, causal_invariance_nonlinear], confounder information (i.e. entire graph) [riskextrap, 15, cocycle] or some proxy thereof in the form of environments [arjovsky2020invariant, ac-vocal, owb-comain-rl, riskextrap]. We, however, do not assume access to any of these and instead synthesize interventions via DA.
Counterfactual DA strategies have been the primary lens for causal analyses of DA [ilse2020selecting, notJustPrettyPictures, causalDA4LLM, mocoda, caiac, causal-matching, contrastive]. These aim for prediction robustness to treatment interventions via DA simulated counterfactuals. [^8] Representing an SEM with exogenous noise distribution conditioned on some variable $Y={\mathbf{y}}$ by $\mathfrak{A}_{Y={\mathbf{y}}}$, the counterfactual SEM $\mathfrak{A}_{Y={\mathbf{y}}};\operatorname{do}\left(X\coloneqq{\mathbf{x}}\right)$ is an intervention $\operatorname{do}\left(X\coloneqq{\mathbf{x}}\right)$ on $\mathfrak{A}_{Y={\mathbf{y}}}$. The resulting counterfactual distribution then captures questions like: “After observing $Y={\mathbf{y}}$, what would have been had $X={\mathbf{x}}$ been true.” As with counterfactual reasoning more broadly, this requires strong assumptions—such as access to the full SEM [notJustPrettyPictures, causalDA4LLM], auxiliary variables [ilse2020selecting, causalDA4LLM, causal-matching, contrastive], or causal graphs [mocoda, caiac]. By contrast, we show that outcome invariance of DA suffices for treatment intervention robustness without invoking counterfactuals. Moreover, prior work has largely overlooked causal effect estimation, often assuming reverse-causal settings where the ATE becomes trivial [ilse2020selecting, causalDA4LLM, notJustPrettyPictures]. Ours is the first framework to study ATE estimation via DA with minimal structural assumptions.
Invariant prediction based methods aim to make predictions based on statistical relationships that remain stable across all domains in $\mathcal{P}$}. A common assumption, for instance, is that $\mathbb{P}_{{Y\mid X}}$ is invariant across $\mathcal{P}$}, with only the marginal $\mathbb{P}_{{X}}$ being allowed to vary. Invariance is also closely linked to causal discovery—following the classic ICM principle [eoci], causal mechanisms remain stable under interventions on inputs [causal-ood-generalizatoin, when-shift-happens]. This connection has inspired approaches that enforce invariance conditions to recover causal structures [icp, causal_invariance_nonlinear]. IV regression can also be viewed as one such method, where the goal is to learn predictors whose residuals are invariant to the instruments [iv_rui, iv_arthur, dual-iv, deep-iv-arthur, specIV]. More broadly, the principle of invariance, whether motivated by causality or otherwise, has proven useful for improving prediction generalization across heterogeneous settings [anchor, arjovsky2020invariant, bo, regularizing-towards-causal-invariance, 7DA, fanny, zz1-droid, rice, krikamol-domain-generalization].
6 Experiments
We began by presenting results in the infinite-sample setting to emphasize that mitigating confounding bias is fundamentally not a sample size issue, i.e., not solvable through traditional regularization alone. In this section, we turn to the finite-sample regime and empirically evaluate the effectiveness of DA in reducing hidden confounding bias. Importantly, we do not use DA for its conventional purpose of augmenting data to improve i.i.d. generalization or reduce estimation variance. Throughout all experiments, we therefore fix the number of samples in the augmented dataset to match that of the original dataset since our focus lies squarely on robust prediction via confounding bias mitigation.
Finding baselines for evaluating our results is however a challenge—the problem of mitigating confounding bias given only observational $(X,Y)$ data and symmetry knowledge via DA is quite underexplored. Nevertheless, for the sake of completeness we make an effort to re-purpose existing methods from domain generalization, invariance learning and causal inference literature to be used as baselines. These methods often require access to additional variables (e.g. IVs, confounders, domains/environments, etc.), and to maintain fairness we will replace these with DA parameters $G$. Such a comparison is conceptually valid since by virtue of being DG methods, they are essentially solving a robust loss of a similar form as in Corollary˜1, giving us meaningful baselines for DA+IVL.
In addition to standard ERM, DA and IV regression, our baselines include DRO [dro], invariant risk minimization (IRM) [arjovsky2020invariant], invariant causal prediction (ICP) [icp], regularization with invariance on causal essential set (RICE) [rice], variance risk extrapolation (V-REx) and minimax risk extrapolation (MM-REx) [riskextrap]. We also include the causal regularization method by Kania and Wit [causal-regularization2] and the $\ell_{1},\ell_{2}$ approaches by Janzing [causal-regularization]. We discretise $G$ if the method accepts only discrete variables. For IVL regression, we select the regularization parameter $\alpha$ in a variety of ways, including vanilla cross validation (CV), level-based CV (LCV) and confounder correction (CC) as described in LABEL:app:implementation-details. Other implementation details are provided in LABEL:app:experiments, and the code to reproduce our results is publicly released at https://github.com/uzairakbar/causal-data-augmentation.
To make CER based evaluation more interpretable for our experiments, we propose the normalization
$\operatorname{nCER}_{\mathfrak{M}}\left(h\right)\coloneqq\frac{\operatorname{CER}_{\mathfrak{M}}\left(h\right)}{\operatorname{CER}_{\mathfrak{M}}\left(h\right)+\operatorname{CER}_{\mathfrak{M}}\left(h_{0}\right)}\in\left[0,1\right],$ $h_{0}\left(\cdot\right)\coloneqq\mathbb{E}^{\mathfrak{M};\operatorname{do}\left(X\right)}\left[Y\right],$
where $h_{0}$ represents the null treatment effect, i.e. when $X$ has no causal influence on $Y$, then $\mathbb{E}^{\mathfrak{M};\operatorname{do}\left(X\right)}\left[Y\;\middle|\;X\right]=\mathbb{E}^{\mathfrak{M};\operatorname{do}\left(X\right)}\left[Y\right]$. The normalized CER (nCER) can be considered a generalization of the metrics used by [causal-regularization, optical-device, dominik-optical-device-causal-reg2] in linear settings and similarly has the interesting property that it is 0 for the ground-truth causal solution $h=f\neq h_{0}$ but $1$ if there is pure confounding for $h\neq f=h_{0}$. Janzing argues in [optical-device, dominik-optical-device-causal-reg2] that using an Euclidean norm instead of the weighted norm in Eq.˜8 is more relevant for causal settings, which also motivates our choice when evaluating results of the simulation and optical-device experiments described below. Conceptually, this is equivalent to evaluation based on the causal risk in Eq.˜3 under the interventional distribution $X^{\prime}\sim{}$